OpenStudy (imsleeziboii):

Simplify ^3 sqrt 432

OpenStudy (imsleeziboii):

@ParthKohli

OpenStudy (anonymous):

can you repeat the question

OpenStudy (imsleeziboii):

simplify \[^{3}\]\[^{3}\sqrt{432}\]

OpenStudy (imsleeziboii):

forget the first ^3

OpenStudy (anonymous):

oh ok do you have a calculator?

OpenStudy (imsleeziboii):

no i dont.

563blackghost (563blackghost):

Is this multiple choice?

OpenStudy (imsleeziboii):

yes

OpenStudy (imsleeziboii):

A.\[^{3}\sqrt{432}\] B.\[2^{3}\sqrt{54}\] C.\[3^{3}\sqrt{16}\] D.\[6^{3}\sqrt{2}\]

OpenStudy (anonymous):

ohhh so then it would be D

563blackghost (563blackghost):

please explain why ^^

OpenStudy (imsleeziboii):

Are you sure?

OpenStudy (anonymous):

cuz \[\sqrt[3]{432} = \sqrt[3]{256 \times 2} = 6 \times \sqrt[3]{2}\]

OpenStudy (anonymous):

you can't simplify 2 any more but the other answers you can

OpenStudy (imsleeziboii):

Can you help me with one more?

OpenStudy (anonymous):

yeah

OpenStudy (imsleeziboii):

Okay

OpenStudy (imsleeziboii):

Which is equivalent to \[^{3}\sqrt{27x ^{6}}\] it has been simplified completely?

OpenStudy (imsleeziboii):

A. 3x^2 B. 3x^3 C. 9x^2 D. 9x^3

OpenStudy (anonymous):

B is the answer

OpenStudy (imsleeziboii):

Thanks

OpenStudy (anonymous):

you take the cube root of 27 and the cube root of x^6 you get 3 * x^3

OpenStudy (imsleeziboii):

Can you help one more time?

OpenStudy (imsleeziboii):

Rewrite the radicand as two factors, one of which is a perfect cube.

OpenStudy (imsleeziboii):

\[3\sqrt{250}\]